# Data-driven Energy Efficiency Modelling in Large-scale Networks: An Expert Knowledge and ML-based Approach

### v1

@@ IV-A Engineering parameters
Parameter
Definition
${\cal RU}$
Set of radio units, ${\cal RU}=\{1,\cdots,r,\cdots,N_{\mathrm{RU}}\}$
$N_{\mathrm{RU}}$
Number of radio units
${\cal C}$
Set of cells, ${\cal C}=\{1,\cdots,i,\cdots,N\}$
$N$
Number of cells
${\cal C}_{r}$
Set of cells operated by radio unit, $r$
$f_{i}$
Carrier frequency of cell, $i$
$B_{i}$
Bandwidth of cell, $i$
$P_{i}^{\rm max}$
Maximum transmit power of cell, $i$
$N_{i}^{\rm DL}$
Number of available DL PRBs in cell, $i$
$N_{i}^{\rm UL}$
Number of available UL PRBs in cell, $i$
@@
Parameter
Definition
$\tau_{i}^{\rm A4}$
A4 handover threshold in cell, $i$
$H_{i}^{\rm A4}$
A4 handover hysteresis in cell, $i$
$O_{i,j}^{\rm freq}$
A4 handover cell-specific frequency offset in cell, $i$, for the inter-frequency neighboring cell, $j$
$O_{i,j}^{\rm cell}$
A4 handover cell-individual offset in cell, $i$, for the inter-frequency neighboring cell, $j$
@@
Parameter
Definition
${\cal C_{\rm C}}$
Set of capacity cells, ${\cal C_{\rm C}}=\{1,\cdots,c,\cdots,N_{\rm C}\}$
$C$
Number of capacity cells
${\cal C_{\rm B}}$
Set of coverage cells, ${\cal C_{\rm B}}=\{1,\cdots,j,\cdots,N_{\rm B}\}$
$B$
Number of coverage cells
$b(c)$
Paired coverage cell of capacity cell, $c$
@@
$\chi^{\rm UE}_{c}$
RRC connected UE threshold in capacity cell, $c$
$\chi^{\rm DL}_{c,b(c)}$
DL PRB threshold in capacity cell, $c$, w.r.t. its paired coverage cell, $b(c)$
$\chi^{\rm UL}_{c,b(c)}$
UL PRB threshold in capacity cell, $c$, w.r.t. its paired coverage cell, $b(c)$
@@
$\Psi^{\rm UE}_{c}$
RRC connected UE threshold in capacity cell, $c$
$\Psi^{\rm DL}_{c,b(c)}$
DL PRB threshold in capacity cell, $c$, w.r.t. its paired coverage cell, $b(c)$
$\Psi^{\rm UL}_{c,b(c)}$
UL PRB threshold in capacity cell, $c$, w.r.t. its paired coverage cell, $b(c)$
@@ IV-B Cell-level KPIs
Parameter
Definition
$U^{\rm UE}_{i,h}$
average number of RRC connected UEs in the cell, $i$, at hour, $h$
$\Delta^{\rm DL}_{i,h}$
average number of DL PRBs used per TTI in cell, $i$, at hour, $h$
$\Delta^{\rm UL}_{i,h}$
average number of UL PRBs used per TTI in cell, $i$, at hour, $h$
$V^{\rm DL}_{i,h}$
number of DL bits successfully transmitted at the radio link control (RLC) layer in cell, $i$, at hour, $h$
$V^{\rm DL-}_{i,h}$
number of DL bits successfully transmitted in the last time slots during which the buffer becomes empty at the RLC layer in cell, $i$, at hour, $h$
$V^{\rm UL}_{i,h}$
number of UL bits successfully received at the RLC layer in cell, $i$, at hour, $h$
$T^{\rm DL}_{i,h}$
amount of time in which cell, $i$, was transmitting DL bits at hour, $h$
$T^{\rm DL-}_{i,h}$
amount of time in which cell, $i$, was transmitting DL bits at hour, $h$, excluding the last time slots during which the DL buffer becomes empty
$T^{\rm UL}_{i,h}$
amount of time in which cell, $i$, was receiving UL bits at hour, $h$
$RN^{\rm DL}_{g,i,h}$
number of samples with the UE rate falling into the $g$-th predefined UE rate range in cell, $i$, at hour, $h$, where $g\in\{1,\cdots,15\}$
@@
Parameter
Definition
$t^{\rm CS}_{i,h}$
duration of carrier shutdown in capacity cell, $c$, at hour $h$
$E^{\rm RU}_{r,h}$
energy consumption of radio unit, $r$, at hour $h$
@@ IV-C UE measurement reports
Parameter
Definition
$T$
Timestamp of measurement
$ID_{u}$
ID of UE, $u$
$S^{\rm ID}_{u}$
ID of the serving cell of UE, $u$
${\cal N}^{\rm ID}_{u}$
Set of IDs of the neighbouring cells of UE, $u$
$M_{u,i}$
RSRP of neighbouring cell, $i\in S^{\rm ID}_{u}\cup{\cal N}^{\rm ID}_{u}$, measured by of UE, $u$
@@ V ML-based radio unit energy consumption model
Understanding and optimizing energy-saving aspects necessitates a thorough grasp of RAN energy consumption modeling. Accurately approximating RAN energy consumption involves summing the energy expended by its radio and base band units. In this section, we outline the framework underpinning our radio unit energy consumption model, centered around ML. Notably, an artificial neural network (ANN) architecture emerged as our choice to model the energy consumption, $E^{\rm RU}_{r,h}$, per radio unit, $r$, at hour, $h$, after rigorous exploration of various ML models due to its robust performance and broad applicability. While the selection process is not presented here for brevity, a comprehensive account can be found in [36] for further insights into this radio unit energy consumption model and our choices.
@@ V-D Architecture of the model
The fundamental architecture chosen for our proposed ANN model is the multilayer perceptron, which comprises several fully connected layers of neurons [38]. The overall structure of the ANN model is illustrated in Fig. 3.
Figure 3: Architecture of the ANN.
In general, the input layer comprised $n_{\rm I}=N_{\mathrm{RU}}+10*C^{\mathrm{MAX}}$ neurons, where $N_{\mathrm{RU}}$ represents the number of distinct radio unit types modeled by the ANN. In our specific data set, $N_{\mathrm{RU}}$ equated to 24, and the number of features considered per radio type was 10, resulting in an input layer of $n_{\mathrm{I}}=84$ neurons. This architecture configuration allowed for effective representation of input data.
@@ V-F Performance of the model
It is important to mention that the training of the ANN model took approximately 75 minutes, encompassing 1086 iterations with a learning rate of 0.001.
@@ VI ML-based DL UE rate model
Unlike energy consumption, which remains relatively consistent for the same radio unit operating under similar conditions, the UE rate is sensitive to the geographical characteristics and channel conditions of a cell’s deployment area. Therefore, two cells with identical configurations and traffic loads but deployed in different locations could exhibit vastly different average and 5%-tile UE rates. To capture these nuances effectively, our experiments showed that using separate models for individual cells is more appropriate than using a universal model for all cells. Furthermore, when pursuing this approach, gradient boosting not only yield improved accuracy, but also reduced complexity in terms of training time compared to an ANN architecture or other alternatives we tested.
@@ VI-A Feature importance analysis
Similar to the approach described in Section V-A, we conducted a feature importance analysis on the collected data sets to determine the key features for estimating average and 5%-tile UE rates. For consistency, we also utilized the same data sets as in Section V-A. However, it is important to note that our cell-level KPIs data sets do not directly provide features corresponding to the average, $R^{\rm avg}_{i,h}$, or 5%-tile, $R^{\rm ce}_{i,h}$, UE rates for each cell, $i$, at hour, $h$:
@@
To provide a comprehensive analysis across all cells, we identified the five most contributing features in each cell based on their SHAP values and ranked them according to their frequency of occurrence. Fig. 6 illustrates how frequently a feature is included in the five most influential features across all cells and UE rate models. This graph underscores the significance of the DL PRB load, DL traffic volume, number of RRC connected UEs, and the fraction of RSRP samples below $-100$ dBm in predicting UE rates.
@@ VI-B Inputs and output of the model
Before constructing our gradient boosting models, we conducted pre-processing on the selected input features as discussed in Section VI-A. An important insight from our data analysis was that cells with high DL PRB load levels tend to have more accurately predicted average and 5%-tile UE rates.
In Fig. 7, we observe the average UE rate of a cell with a low number of UEs and consistently low loads, while Fig. 8 depicts a cell with higher UE count and larger average loads. The former graph highlights the considerable variance in average UE rates when $\frac{\Delta^{\rm DL}_{i,h}}{N_{i}^{\rm DL}}<0.1$, which impacts the accuracy of UE rate estimation using available data.
In intuitive terms, when the DL PRB load, $\frac{\Delta^{\rm DL}_{i,h}}{N_{i}^{\rm DL}}$, is low, the UEs sharing the transmission slot might have limited payload, potentially leading to underutilized bandwidth. Consequently, the average UE rate, $R^{\rm avg}_{i,h}$, becomes highly dependent on the specific traffic characteristics, such as file sizes. For instance, if there are only enough bits to occupy 25 % or 75 % of the transmission slot’s bandwidth, the resulting UE rate would be 25 % or 75 % of the cell’s maximum capacity. As accurately predicting traffic nature and file sizes based on available data is challenging, we decided to model average and 5%-tile UE rates using our gradient boosting approach, focusing solely on input data corresponding to DL PRB load smaller or equal than 0.1, i.e. $\frac{\Delta^{\rm DL}_{i,h}}{N_{i}^{\rm DL}}\geq 0.1$. For lower loads, they were modeled using a random variable with its probability mass function (PMF) learned from the data.
@@ VI-D Performance of the model
To evaluate performance, we conducted a comparison between the estimated average and 5%-tile percentile UE rates using both the ideal baseline and ABM-friendly estimators, against the UE rates obtained from measurements in the test set. This assessment employed the formulation introduced in Section VI-A. For this comparison, we employed the MAE and MAPE metrics. Additionally, we employed the mean estimator as a further benchmark, which characterizes the DL UE rate by computing the mean values of the related data observed in the training set.
Regarding complexity, it is worth noting that our gradient boosting model required less than 1 minute per cell for training the ideal baseline or the ABM-friendly estimator. The hyperparameters of the gradient boosting model were configured as follows: 500 estimators, a maximum depth of 4, and a learning rate of 0.01.
The results, presented in Table V, reveal that, as anticipated, the ideal baseline estimator exhibits the highest performance when evaluating the average and 5%-tile percentile UE rates. Specifically, it achieves a MAE of 2.24 Mbps and 0.14 Mbps, alongside a MAPE of 10.60 % and 27.94 % for the average and 5%-tile percentile UE rates, respectively. The 5%-tile percentile UE rates, characterized by lower values than the average, incurs larger MAPE values.
In contrast, the mean estimator demonstrates the weakest performance, yielding a MAE of 4.27 Mbps and 0.14 Mbps, along with a MAPE of 21.52 % and 35.9 %, for the average and 5%-tile percentile UE rates, respectively. This represents a 2$\times$ and a 28 % increase in MAPE when estimating the mean and the 5%-tile UE rate, respectively.
@@ VII-A Problem statement
Outputs of our expert-based, white-box model: We denote the set of stochastic outputs of our expert-based, white-box model as ${\cal X}^{\rm OUT}_{i,h^{\prime}}$, wherein this set includes the predicted distributions, $\hat{f}_{U^{\rm UE}_{i,h^{\prime}}}$, $\hat{f}_{\Delta^{\rm DL}_{i,h^{\prime}}}$, and $\hat{f}_{t^{\rm CS}_{i,h^{\prime}}}$, of the predicted number $\hat{U}^{\rm UE}_{i,h^{\prime}}$ of RRC connected UEs, the predicted number, $\hat{\Delta}^{\rm DL}_{i,h^{\prime}}$, of utilized DL PRBs, and the predicted duration, $\hat{t}^{\rm CS}_{i,h^{\prime}}$, of carrier shutdown, for each cell, $i$, and hour, hsuperscriptnormal-$h^{\prime}$, respectively.
@@ VII-B Dealing with stochasticity - Monte Carlo method
Outputs of each Monte Carlo run, r: Let ${\cal X}^{\rm MC-OUT}_{i,h^{\prime},r}$ symbolize the set of outputs for each Monte Carlo run, r, wherein this set encompasses the predicted number, $\hat{U}^{\rm UE}_{i,h^{\prime},r}$, of RRC connected UEs, the predicted number, $\hat{\Delta}^{\rm DL}_{i,h^{\prime},r}$, of used DL PRBs, and the predicted duration, $\hat{t}^{\rm CS}_{i,h^{\prime},r}$, of the carrier shutdown for each cell, $i$, hour, hsuperscriptnormal-$h^{\prime}$, and Monte Carlo run, r.

### v2

@@ IV-A Engineering parameters
Parameter Definition
${\cal RU}$ Set of radio units, ${\cal RU}=\{1,\cdots,r,\cdots,N_{\mathrm{RU}}\}$
$N_{\mathrm{RU}}$ Number of radio units
${\cal C}$ Set of cells, ${\cal C}=\{1,\cdots,i,\cdots,N\}$
$N$ Number of cells
${\cal C}_{r}$ Set of cells operated by radio unit, $r$
$f_{i}$ Carrier frequency of cell, $i$
$B_{i}$ Bandwidth of cell, $i$
$P_{i}^{\rm max}$ Maximum transmit power of cell, $i$
$N_{i}^{\rm DL}$ Number of available DL PRBs in cell, $i$
$N_{i}^{\rm UL}$ Number of available UL PRBs in cell, $i$
@@
Parameter Definition
$\tau_{i}^{\rm A4}$ A4 handover threshold in cell, $i$
$H_{i}^{\rm A4}$ A4 handover hysteresis in cell, $i$
$O_{i,j}^{\rm freq}$ A4 handover cell-specific frequency offset in cell, $i$, for the inter-frequency neighboring cell, $j$
$O_{i,j}^{\rm cell}$ A4 handover cell-individual offset in cell, $i$, for the inter-frequency neighboring cell, $j$
@@
Parameter Definition
${\cal C_{\rm C}}$ Set of capacity cells, ${\cal C_{\rm C}}=\{1,\cdots,c,\cdots,N_{\rm C}\}$
$C$ Number of capacity cells
${\cal C_{\rm B}}$ Set of coverage cells, ${\cal C_{\rm B}}=\{1,\cdots,j,\cdots,N_{\rm B}\}$
$B$ Number of coverage cells
$b(c)$ Paired coverage cell of capacity cell, $c$
@@
$\chi^{\rm UE}_{c}$ RRC connected UE threshold in capacity cell, $c$
$\chi^{\rm DL}_{c,b(c)}$ DL PRB threshold in capacity cell, $c$, w.r.t. its paired coverage cell, $b(c)$
$\chi^{\rm UL}_{c,b(c)}$ UL PRB threshold in capacity cell, $c$, w.r.t. its paired coverage cell, $b(c)$
@@
$\Psi^{\rm UE}_{c}$ RRC connected UE threshold in capacity cell, $c$
$\Psi^{\rm DL}_{c,b(c)}$ DL PRB threshold in capacity cell, $c$, w.r.t. its paired coverage cell, $b(c)$
$\Psi^{\rm UL}_{c,b(c)}$ UL PRB threshold in capacity cell, $c$, w.r.t. its paired coverage cell, $b(c)$
@@ IV-B Cell-level KPIs
Parameter Definition
$U^{\rm UE}_{i,h}$ average number of RRC connected UEs in the cell, $i$, at hour, $h$
$\Delta^{\rm DL}_{i,h}$ average number of DL PRBs used per TTI in cell, $i$, at hour, $h$
$\Delta^{\rm UL}_{i,h}$ average number of UL PRBs used per TTI in cell, $i$, at hour, $h$
$V^{\rm DL}_{i,h}$ number of DL bits successfully transmitted at the radio link control (RLC) layer in cell, $i$, at hour, $h$
$V^{\rm DL-}_{i,h}$ number of DL bits successfully transmitted in the last time slots during which the buffer becomes empty at the RLC layer in cell, $i$, at hour, $h$
$V^{\rm UL}_{i,h}$ number of UL bits successfully received at the RLC layer in cell, $i$, at hour, $h$
$T^{\rm DL}_{i,h}$ amount of time in which cell, $i$, was transmitting DL bits at hour, $h$
$T^{\rm DL-}_{i,h}$ amount of time in which cell, $i$, was transmitting DL bits at hour, $h$, excluding the last time slots during which the DL buffer becomes empty
$T^{\rm UL}_{i,h}$ amount of time in which cell, $i$, was receiving UL bits at hour, $h$
$RN^{\rm DL}_{g,i,h}$ number of samples with the UE rate falling into the $g$-th predefined UE rate range in cell, $i$, at hour, $h$, where $g\in\{1,\cdots,15\}$
@@
Parameter Definition
$t^{\rm CS}_{i,h}$ duration of carrier shutdown in capacity cell, $c$, at hour $h$
$E^{\rm RU}_{r,h}$ energy consumption of radio unit, $r$, at hour $h$
@@ IV-C UE measurement reports
Parameter Definition
$T$ Timestamp of measurement
$ID_{u}$ ID of UE, $u$
$S^{\rm ID}_{u}$ ID of the serving cell of UE, $u$
${\cal N}^{\rm ID}_{u}$ Set of IDs of the neighbouring cells of UE, $u$
$M_{u,i}$ RSRP of neighbouring cell, $i\in S^{\rm ID}_{u}\cup{\cal N}^{\rm ID}_{u}$, measured by of UE, $u$
@@ V ML-based radio unit energy consumption model
Understanding and optimizing energy-saving aspects necessitates a thorough grasp of RAN energy consumption modeling. Accurately approximating RAN energy consumption involves summing the energy expended by its radio and base band units. In this section, we outline the framework underpinning our radio unit energy consumption model, centered around ML. Notably, an ANN architecture emerged as our choice to model the energy consumption, $E^{\rm RU}_{r,h}$, per radio unit, $r$, at hour, $h$, after rigorous exploration of various ML models due to its robust performance and broad applicability. While the selection process is not presented here for brevity, a comprehensive account can be found in [65] for further insights into this radio unit energy consumption model and our choices.
@@ V-D Architecture of the model
The fundamental architecture chosen for our proposed ANN model is the multilayer perceptron, which comprises several fully connected layers of neurons [67]. Fig. 3 offers a comprehensive visualization of the overall ANN employed for estimating power consumption in this paper.
Figure 3: Structure of our proposed ANN, highlighting the selected input variables, the organization and interconnection of the two hidden layers, and the configuration of the two output neurons which represent the mean and standard deviation of the power consumption estimates.
Following from the figure, the input layer comprised $n_{\rm I}=N_{\mathrm{RU}}+10*C^{\mathrm{MAX}}$ neurons, where $N_{\mathrm{RU}}$ represents the number of distinct radio unit types modeled by the ANN. In our specific data set, $N_{\mathrm{RU}}$ equated to 24, and the number of features considered per radio type was 10, resulting in an input layer of $n_{\mathrm{I}}=84$ neurons. This architecture configuration allowed for effective representation of input data.
@@ V-F Performance of the model
It is important to mention that the training of the ANN model took approximately 75 minutes in a machine powered by a Intel®  CoreTM i7-9700 CPU @ 3.00 GHz with 32 GB of Random Access Memory (RAM), encompassing 1086 iterations with a learning rate of 0.001.
@@ VI ML-based DL UE rate model
Unlike energy consumption, which remains relatively consistent for the same radio unit operating under similar conditions, the UE rate is sensitive to the geographical characteristics and channel conditions of a cell’s deployment area. Therefore, two cells with identical configurations and traffic loads but deployed in different locations could exhibit vastly different average and 5%-tile UE throughput. To capture these nuances effectively, our experiments showed that using separate models for individual cells is more appropriate than using a universal model for all cells. Furthermore, when pursuing this approach, gradient boosting not only yield improved accuracy, but also reduced complexity in terms of training time compared to an ANN architecture or other alternatives we tested.
@@ VI-A Feature importance analysis
Similar to the approach described in Section V-A, we conducted a feature importance analysis on the collected data sets to determine the key features for estimating average and 5%-tile UE throughput. For consistency, we also utilized the same data sets as in Section V-A. However, it is important to note that our cell-level KPIs data sets do not directly provide features corresponding to the average, $R^{\rm avg}_{i,h}$, or 5%-tile, $R^{\rm ce}_{i,h}$, UE throughput for each cell, $i$, at hour, $h$:
@@
To provide a comprehensive analysis across all cells, we identified the five most contributing features in each cell based on their SHAP values and ranked them according to their frequency of occurrence. Fig. 6 illustrates how frequently a feature is included in the five most influential features across all cells and UE rate models. This graph underscores the significance of the DL PRB load, DL traffic volume, number of RRC connected UEs, and the fraction of RSRP samples below $-100$ dBm in predicting UE throughput.
@@ VI-B Inputs and output of the model
Before constructing our gradient boosting models, we conducted pre-processing on the selected input features as discussed in Section VI-A. An important insight from our data analysis was that cells with high DL PRB load levels tend to have more accurately predicted average and 5%-tile UE throughput.
In Fig. 7, we observe the average UE rate of a cell with a low number of UEs and consistently low loads, while Fig. 8 depicts a cell with higher UE count and larger average loads. The former graph highlights the considerable variance in average UE throughput when $\frac{\Delta^{\rm DL}_{i,h}}{N_{i}^{\rm DL}}<0.1$, which impacts the accuracy of UE rate estimation using available data.
In intuitive terms, when the DL PRB load, $\frac{\Delta^{\rm DL}_{i,h}}{N_{i}^{\rm DL}}$, is low, the UEs sharing the transmission slot might have limited payload, potentially leading to underutilized bandwidth. Consequently, the average UE rate, $R^{\rm avg}_{i,h}$, becomes highly dependent on the specific traffic characteristics, such as file sizes. For instance, if there are only enough bits to occupy 25 % or 75 % of the transmission slot’s bandwidth, the resulting UE rate would be 25 % or 75 % of the cell’s maximum capacity. As accurately predicting traffic nature and file sizes based on available data is challenging, we decided to model average and 5%-tile UE throughput using our gradient boosting approach, focusing solely on input data corresponding to DL PRB load smaller or equal than 0.1, i.e. $\frac{\Delta^{\rm DL}_{i,h}}{N_{i}^{\rm DL}}\geq 0.1$. For lower loads, they were modeled using a random variable with its probability mass function (PMF) learned from the data.
@@ VI-D Performance of the model
To evaluate performance, we conducted a comparison between the estimated average and 5%-tile percentile UE throughput using both the ideal baseline and ABM-friendly estimators, against the UE throughput obtained from measurements in the test set. This assessment employed the formulation introduced in Section VI-A. For this comparison, we employed the MAE and MAPE metrics. Additionally, we employed the mean estimator as a further benchmark, which characterizes the DL UE rate by computing the mean values of the related data observed in the training set.
Regarding complexity, it is worth noting that our gradient boosting model required less than 1 minute per cell for training the ideal baseline or the ABM-friendly estimator in a machine powered by a Intel®  CoreTM i7-9700 CPU @ 3.00 GHz with 32 GB of Random Access Memory. The hyperparameters of the gradient boosting model were configured as follows: 500 estimators, a maximum depth of 4, and a learning rate of 0.01.
@@
The results, presented in Table V, reveal that, as anticipated, the ideal baseline estimator exhibits the highest performance when evaluating the average and 5%-tile percentile UE throughput. Specifically, it achieves a MAE of 2.24 Mbps and 0.14 Mbps, alongside a MAPE of 10.60 % and 27.94 % for the average and 5%-tile percentile UE throughput, respectively. The 5%-tile percentile UE throughput, characterized by lower values than the average, incurs larger MAPE values.
In contrast, the mean estimator demonstrates the weakest performance, yielding a MAE of 4.27 Mbps and 0.14 Mbps, along with a MAPE of 21.52 % and 35.9 %, for the average and 5%-tile percentile UE throughput, respectively. This represents a 2$\times$ and a 28 % increase in MAPE when estimating the mean and the 5%-tile UE rate, respectively.
@@ VII-A Problem statement
Outputs of our expert-based, white-box model: We denote the set of stochastic outputs of our expert-based, white-box model as ${\cal X}^{\rm OUT}_{i,h^{\prime}}$, wherein this set includes the predicted distributions, $\hat{f}_{U^{\rm UE}_{i,h^{\prime}}}$, $\hat{f}_{\Delta^{\rm DL}_{i,h^{\prime}}}$, and $\hat{f}_{t^{\rm CS}_{i,h^{\prime}}}$, of the predicted number $\hat{U}^{\rm UE}_{i,h^{\prime}}$ of RRC connected UEs, the predicted number, $\hat{\Delta}^{\rm DL}_{i,h^{\prime}}$, of utilized DL PRBs, and the predicted duration, $\hat{t}^{\rm CS}_{i,h^{\prime}}$, of carrier shutdown, for each cell, $i$, and hour, $h^{\prime}$, respectively.
@@ VII-B Dealing with stochasticity - Monte Carlo method
Outputs of each Monte Carlo run, r: Let ${\cal X}^{\rm MC-OUT}_{i,h^{\prime},r}$ symbolize the set of outputs for each Monte Carlo run, r, wherein this set encompasses the predicted number, $\hat{U}^{\rm UE}_{i,h^{\prime},r}$, of RRC connected UEs, the predicted number, $\hat{\Delta}^{\rm DL}_{i,h^{\prime},r}$, of used DL PRBs, and the predicted duration, $\hat{t}^{\rm CS}_{i,h^{\prime},r}$, of the carrier shutdown for each cell, $i$, hour, $h^{\prime}$, and Monte Carlo run, r.